OpenStudy (anonymous):

may someone help me with the least common denominator of 8/5/3

OpenStudy (cwrw238):

8 = 2*2*2 5 3 we pair numbers : there is one pair of 2's - so we only use one of these: LCM = 2*2*3*5 = 60

OpenStudy (anonymous):

I don't think eight goes into 60. By my calculation, the LCM is 120=2^3*3*5

OpenStudy (cwrw238):

yes - sorry - i got confused - the pairs should be on different numbers u r right